Write the body of your question here and enumerate the parts as indicated below:

(a) First part

(b) Second part. This question has subparts, which should be enumerated as below. (i) First subpart (ii) Second subpart (iii) Third subpart

2. Here comes another question...

(a) First part

(b) Second part. This question has subparts, which should be enumerated as below. (i) First subpart (ii) Second subpart (iii) Third subpart

JAST  1  Turn over